# Multi-level projection with exponential parallel speedup; Application to sparse auto-encoders neural networks

### v1

@@ Abstract
The $\ell_{1,\infty}$ norm is an efficient structured projection but the complexity of the best algorithm is unfortunately $\mathcal{O}\big{(}nm\log(nm)\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$. In this paper, we propose a new bi-level projection method for which we show that the time complexity for the $\ell_{1,\infty}$ norm is only $\mathcal{O}\big{(}nm\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$, and $\mathcal{O}\big{(}n+m\big{)}$ with full parallel power. We generalize our method to tensors and we propose a new multi-level projection, having an induced decomposition that yields a linear parallel speedup up to an exponential speedup factor, resulting in a time complexity lower-bounded by the sum of the dimensions. Experiments show that our bi-level $\ell_{1,\infty}$ projection is $2.5$ times faster than the actual fastest algorithm provided by Chu et. al. while providing same accuracy and better sparsity in neural networks applications.
@@ 1.2 Contributions
In this paper we focus on the computational issue for the $\ell_{1,\infty}$ norm. First, in order to cope with this issue, we propose a new bi-level projection method. The main motivation for this work is the direct independent splitting made by the bi-level optimization which take into account the structured sparsity requirement. Second, we generalize the bi-level projection to multi-level, and shows its theoretical guaranties, such as an exponential parallel speedup. The paper is organized as follows. First we present our general bi-level projection framework using our splitting approach 3. Then, we provide in section 4 the application to the bi-level $\ell_{1,\infty}$ projection. In section 5, we apply our bi-level framework to well known constraints such as $\ell_{1,1}$ constraint and to any combination of p, q norm such as the group-Lasso and the exclusive LASSO. We extend our bi-level projection method to tensors using multi-level projections in section 6. In Section 7, we finally compare different projection methods experimentally. Our experimental section is split in two parts. First, we provide an empirical analysis of the projection algorithms onto the bi-level projection $\ell_{1,\infty}$ ball. This section shows the benefit of the proposed method, especially in the context of sparsity. Second, we apply our framework to the diagnosis (or classification) using a supervised autoencoder on a synthetic dataset and a biological dataset.
@@ 2 Definitions and Notations
In this paper we use the following notations:lowercase Greek symbol for scalars, scalar i,j,c,m,n are indices of vectors and matrices, lowercase for vectors, capital for matrices, and calligraphic for tensors. For every $p,q\in\mathbb{R}$, the $\ell_{p,q}$ norm of a real matrix $X=[x_{1}\ \cdots\ x_{m}]\in\mathbb{R}^{n\times m}$ with columns $x_{j}$ and elements $X_{i,j}$ is given by:
@@ 3 A new bi-level projection formulation
Processing the projection of matrices or other mathematical objects, such as tensors, is becoming harder and harder as dedicated algorithms satisfying equation (4) must be defined for each possible combination of $p$ and $q$. of equation (4) and to reformulate the projection as a bi-level problem. In this reformulated problem, the $\ell_{p,q}$ projection will be split into an aggregation using the $q$ norm followed by a simpler and well defined projection onto the $p$ norm of the aggregated vector. Let $v_{q}=({\|{Y_{1}}\|}_{q},\dots,{\|{Y_{n}}\|}_{q})$ be the vector composed of the $q$ norms of the columns of a matrix.
@@
$BP^{p,q}_{\eta}(Y)=\{X|\forall i,x_{i}=\arg\min\limits_{x_{i}\in\mathcal{B}_{\hat{u}_{i}}^{q}}{\|{X_{i}-Y_{i}}\|}_{2}$ (5)
@@
Then, the second part of the bi-level optimization problem, once vector $\hat{u}$ is known is given by:
$x_{i}=\min\limits_{x_{i}\in\mathcal{B}_{\hat{u}_{i}}^{q}}{\|{x_{i}-y_{i}}\|}_{2}$ (8)
For each column $i$ of the original matrix. Here, for each $x_{i}$, an independent optimization problem is defined. Each of them is optimally solved by definition of the projection on the $\ell_{q}$ ball:
$x_{i}\leftarrow P^{q}_{u_{i}}(y_{i})\forall i\in 1,\dots,n$ (9)
@@
  $u\leftarrow P^{p}_{\eta}(({\|{y_{1}}\|}_{q},{\|{y_{j}}\|}_{q},\dots,{\|{y_{n}}\|}_{q}))$
  for $i\in[1,\dots,n]$ do
     $x_{i}\leftarrow P^{q}_{u_{i}}(y_{i})$
@@ 4.1 A new bi-level projection
Given a radius $\eta\geq 0$, the goal is to project $Y$ onto the $\ell_{1,\infty}$ norm ball of radius $C$, denoted by
@@
In this paper we propose the following alternative new bi-level method. Recall that ${v_{\infty}}=({\|{Y_{1}}\|}_{\infty},\dots,{\|{Y_{n}}\|}_{\infty})$ is the vector composed of the infinity norms of the columns of matrix $Y$. Let the infinity norm projection $P^{\infty}_{u_{i}}(y)=(\min(y_{i},u_{i}),\forall y_{i}\in y)$. The bi-level $\ell_{1,\infty}$ projection optimization problem is defined by:
$BP^{1,\infty}_{\eta}(Y)=\{X|\forall i,X_{i}=\arg\min\limits_{X_{i}\in\mathcal{B}_{\hat{u}_{i}}^{\infty}}{\|{X_{i}-Y_{i}}\|}_{2}$ (13)
@@
  $u\leftarrow P^{1}_{\eta}(({\|{y_{1}}\|}_{\infty},{\|{y_{j}}\|}_{\infty},\dots,{\|{y_{n}}\|}_{\infty}))$
  for $i\in[1,\dots,n]$ do
     $x_{i}\leftarrow P^{\infty}_{u_{i}}(y_{i})$
@@ 4.2 Computational complexity
Synthetic (bi-level/usual) $\ell_{1,2}$ $\ell_{1,1}$ bi-level $\ell_{1,1}$ $\ell_{1,\infty}$ bi-level $\ell_{1,\infty}$
@@
5 Extension to other norms
5.1 Extension to the $\ell_{1,1}$ bi-level projection
The $\ell_{1}$ ball is famous for being robust and yielding sparsity. Nevertheless, its extension to matrix, the $\ell_{1,1}$ does not yield structured sparsity. We propose to define the $\ell_{1,1}$ bi-level optimization problem:
$BP^{1,1}_{\eta}(Y)=\{X|\forall i,X_{i}=\arg\min\limits_{X_{i}\in\mathcal{B}_{\hat{u}_{i}}^{1}}{\|{X_{i}-Y_{i}}\|}_{2}$ (14)
@@ 5.1 Extension to the $\ell_{1,1}$ bi-level projection
  $u\leftarrow P^{1}_{\eta}(({\|{y_{1}}\|}_{1},\dots,{\|{y_{n}}\|}_{1}))$
  for $i\in[1,\dots,n]$ do
     $x_{i}\leftarrow P_{u_{i}}^{1}(y_{i})$
@@
5.2 Extension to bi-level $\ell_{1,2}$ and $\ell_{2,1}$ projections
Group LASSO and Exclusive LASSO are well-known algorithm with impressive properties for grouping variables and solving inverse problems [16, 31, 32, 33]. For example, the $\ell_{1,2}$ constraint has been proposed in the context of Group Lasso regularization [34], where the main idea was to enforce parameters of different classes to share common features. We propose to consider their bi-level counter-parts (i.e. $\ell_{1,2}$ and $\ell_{2,1}$). Let the $\ell_{1,2}$ bi-level optimization problem be:
$BP^{1,2}_{\eta}(Y)=\{X|\forall i,X_{i}=\arg\min\limits_{X_{i}\in\mathcal{B}_{\hat{u}_{i}}^{2}}{\|{X_{i}-Y_{i}}\|}_{2}$ (15)
@@ 5.2 Extension to bi-level $\ell_{1,2}$ and $\ell_{2,1}$ projections
First, the bi-level projection algorithms for $\ell_{1,2}$ is given by algorithm 4.
@@
  $u\leftarrow P^{1}_{\eta}(({\|{Y_{2}}\|}_{2},\dots,{\|{Y_{n}}\|}_{2}))$
  for $i\in[1,\dots,n]$ do
     $x_{i}\leftarrow P^{2}_{u_{i}}(y_{i})\forall i\in 1,\dots,n$
@@
It is interesting to note that the bi-level $\ell_{1,2}$ projection and usual Group LASSO formulation are different. The usual Group LASSO formulation is provided in Appendix Section A. For the bi-level $\ell_{2,1}$ projection the algorithm is given in appendix. This second algorithm, and the many others that can be defined using our framework, is considered out of the topic of this paper as we focus on structured sparsity for the columns for our matrix norm experiments. Nevertheless, they can be defined and have the same advantages as any assignment of $p$ and $q$.
@@ 6.1 Tri-level Projection Algorithms
Matrices are not the only mathematical objects that are used in neural networks. Images, for examples, are often represented as order 3 tensors in $\mathbb{R}^{c,n,m}$ with $c$ representing the channels of the image. Moreover, the current development of deep-learning framework for image compression are leaning toward tensor efficient regularization methods [35, 36], and not only vectors or matrices methods. For example, the new image compression standard JPEG AI [37] uses tensor representation of images in the latent space of an autoencoder. That is the reason why we propose to generalize the bi-level projection, first to tri-level, then to multi-level. We consider tensors as multi-dimensional arrays with subscripts indexing. First, consider order 3 tensors of the form $\mathcal{Y}\in\mathbb{R}^{c,n,m}$. For example, $\mathcal{Y}\in\mathbb{R}^{3,m,n}$ is used to represent an image of width $m\times n$ and having 3 channels. We define by $\forall i$ $\in<c,n,m>$ all the tuples of the Cartesian product of the indices. For example $\forall i$ <3,256,256$\in$<3,256,256>∈ < 3 , 256 , 256 > $\equiv\forall i\in((1,1,1),(1,1,2),\dots,(3,256,256))$. Given a tensor $\mathcal{Y}\in\mathbb{R}^{c,n,m}$, we denote by $V_{q}=({\|{\mathcal{Y}_{1,1}}\|}_{q},\dots,{\|{\mathcal{Y}_{n,m}}\|}_{q})$ the matrix of aggregation of the channels by the norm $q$.
@@ 7.3.2 Experimental accuracy results
Figure 5 shows the impact of the radius ($\eta$) on synthetic and lung dataset using by the bilevel projection.
Figure 6: Sparsity as a function of the radius parameter $\eta$: Left Synthetic data, Right Lung dataset
It can be seen that tuning the projection radius, and thus the sparsity, is necessary to improve the accuracy for synthetic data. From Table 2 on synthetic dataset and Table 3 on Lung dataset it can be seen that the best accuracy is obtained for $\eta=0.75$ for $\ell_{1,\infty}$ and for $\eta=1.$ for the bilevel $\ell_{1,\infty}$ projection, maximum accuracy of both method are similar. However sparsity and computation time are better for bilevel $\ell_{1,\infty}$ than for regular $\ell_{1,\infty}$.
@@
Synthetic Baseline $\ell_{1,2}$ bi-level $\ell_{1,1}$
Best Radius 75 75
[…]
Table 4: Synthetic dataset. SiLU activation, Accuracy versus sparsity : comparison of $\ell_{1,2}$ and bi-level $\ell_{1,1}$.
The table 4 on synthetic dataset and the Table 5 on Lung dataset show that the projection bi-level $\ell_{1,1}$ and $\ell_{1,2}$ have almost the same accuracy performance as the projection bi-level $\ell_{1,\infty}$ but the sparsity is much lower.
Lung Baseline $\ell_{1,2}$ bi-level $\ell_{1,1}$
Best Radius 75 200
[…]
Table 5: Lung dataset. SiLU activation, Accuracy versus sparsity : comparison of $\ell_{1,2}$ and bi-level $\ell_{1,1}$
So bilevel $\ell_{1,\infty}$ outperforms all other projections in terms of computation cost and sparsity.
@@ 8 Conclusion
Appendix A Group LASSO and Exclusive LASSO
Let first recall the definition of the usual $\ell_{1,2}$ Group LASSO projection be:
$P^{1,2}_{\eta}(Y)=\arg\min\limits{\|{X-Y}\|}_{2}\text{ s.t. }\|X\|_{1,2}\leq\eta.$ (19)
Note that the formulation of the Group-LASSO is different from our bi-level approach, and their optimal points are in the general case different. The classical Group-Lasso algorithm is based on Block coordinate descent (BCD) [20] which requires high computational cost. Recent paper, provided algorithms for fast Group LASSO, yet the algorithm is quite complicated [21].
The exclusive LASSO was first introduced in [31, 32, 33]. The main idea is that if one feature in a class is selected (large weight), the method tends to assign small weights to the other features in the same class. Using our bi-level framework, we obtain the following bi-level projection algorithms.
Algorithm 7 Bi-level $\ell_{2,1}$ projection.
  Input: $Y,\eta$
  $t\leftarrow P^{2}_{\eta}(({\|{Y_{2}}\|}_{1},\dots,{\|{Y_{n}}\|}_{1}))$
  for $i\in[1,\dots,n]$ do
     $X_{i}\leftarrow P^{1}_{t_{i}}(Y_{i})\forall i\in 1,\dots,n$
  end for
  Output: $X$
Appendix B Double descent algorithm
Originally proposed as follows: after training a network, set all weights smaller than a given threshold to zero, rewind the rest of the weights to their initial configuration, and then retrain the network from this starting configuration while keeping the zero weights frozen (untrained). A possible implementation of the double descent algorithm is given in Algorithm 8.
Algorithm 8 Double-descent algorithm.
1:  Input: $W*,\gamma,\eta$
2:  for $n=1,\dots,N(epochs)$ do
3:     $V\leftarrow f(W,\gamma,\nabla\psi(W))$
4:  end for
5:  $W:=BP^{p,q}_{\eta}(V)$
6:  $(M_{0})_{ij}\leftarrow\mathbbm{1}_{x\neq 0}(w_{ij})$
7:  for $n=1,\dots,N(epoch)$ do
8:     $W\leftarrow f(W,\gamma,\nabla\phi(W,M_{0}))$
9:  end for
10:  Output: $W$
In line 3 the first descent is made. Then in line 5 the projection is made and line 6 extract the mask of non-zero entries in the projected matrix. Finally line 8 is the second descent. Note that low values of $\eta$ imply high sparsity of the network.
Appendix C Iterative form of the multi-level projection

### v2

# Multi-level projection with exponential parallel speedup; Application to sparse neural networks

@@ Abstract
The $\ell_{1,\infty}$ norm is an efficient structured projection but the complexity of the best algorithm is unfortunately $\mathcal{O}\big{(}nm\log(nm)\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$. In this paper, we propose a new bi-level projection method for which we show that the time complexity for the $\ell_{1,\infty}$ norm is only $\mathcal{O}\big{(}nm\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$, and $\mathcal{O}\big{(}n+m\big{)}$ with full parallel power. We generalize our method to tensors and we propose a new multi-level projection, having an induced decomposition that yields a linear parallel speedup up to an exponential speedup factor, resulting in a time complexity lower-bounded by the sum of the dimensions, instead of the product of the dimensions. we provide a large base of implementation of our framework for bi-level and tri-level (matrices and tensors) for various norms and provides also the parallel implementation. Experiments show that our projection is $2$ times faster than the actual fastest Euclidean algorithms while providing same accuracy and better sparsity in neural networks applications.
@@ 1.2 Contributions
In this paper we focus on the computational issue for the $\ell_{1,\infty}$ norm. First, in order to cope with this issue, we propose a new bi-level projection method. The main motivation for this work is the direct independent splitting made by the bi-level optimization which take into account the structured sparsity requirement. Second, we generalize the bi-level projection to multi-level, and shows its theoretical guaranties, such as an exponential parallel speedup. The paper is organized as follows. First we present our general bi-level projection framework using our splitting approach 3. Then, we provide in section 4 the application to the bi-level $\ell_{1,\infty}$ projection. In section 4.3, we apply our bi-level framework to well known constraints such as $\ell_{1,1}$ constraint and to any combination of p, q norm. We extend our bi-level projection method to tensors using multi-level projections in section 5. In Section 6, we finally compare different projection methods experimentally. Our experimental section is split in two parts. First, we provide an empirical analysis of the projection algorithms onto the bi-level projection $\ell_{1,\infty}$ ball. This section shows the benefit of the proposed method, especially in the context of sparsity. Second, we apply our framework to the diagnosis (or classification) using a neural network (Autoencoder) on a synthetic dataset and a biological dataset.
@@ 2 Definitions and Notations
In this paper we use the following notations:lowercase Greek symbol for scalars, scalar i,j,c,m,n are indices of vectors and matrices, lowercase for vectors, capital for matrices, and calligraphic for tensors. For every $p,q\in 1,2,\cdot,\infty$, the $\ell_{p,q}$ norm of a real matrix $X=[x_{1}\ \cdots\ x_{m}]\in\mathbb{R}^{n\times m}$ with columns $x_{j}$ and elements $X_{i,j}$ is given by:
@@ 3 A new bi-level projection formulation
Processing the projection of matrices or other mathematical objects, such as tensors, is becoming harder and harder as dedicated algorithms satisfying equation (4) must be defined for each possible combination of $p$ and $q$. In this section we relax equation (4) and to reformulate the projection as a bi-level problem. In this reformulated problem, the $\ell_{p,q}$ projection will be split into an aggregation using the $q$ norm followed by a simpler and well defined projection onto the $p$ norm of the aggregated vector. Let $v_{q}=({\|{y_{1}}\|}_{q},\dots,{\|{y_{m}}\|}_{q})$ be the raw vector composed of the $q$ norms of the columns of the matrix Y.
@@
$BP^{p,q}_{\eta}(Y)=\{X|\forall j,x_{j}=\arg\min\limits_{x_{j}\in\mathcal{B}_{\hat{u}_{j}}^{q}}{\|{x_{j}-y_{j}}\|}_{2}$ (5)
@@
Then, the second part of the bi-level optimization problem, once vector $\hat{u}$ is known, for each column $j$ of the original matrix $x_{j}$ an independent optimization problem is defined:
$x_{j}=\min\limits_{x_{j}\in\mathcal{B}_{\hat{u}_{j}}^{q}}{\|{x_{j}-y_{j}}\|}_{2}$ (8)
Each of column $x_{j}$ is optimally solved by definition of the projection on the $\ell_{q}$ ball:
$x_{j}\leftarrow P^{q}_{u_{j}}(y_{j})\forall j\in 1,\dots,m$ (9)
@@
  $u\leftarrow P^{p}_{\eta}(({\|{y_{1}}\|}_{q},{\|{y_{j}}\|}_{q},\dots,{\|{y_{m}}\|}_{q}))$
  for $j\in[1,\dots,m]$ do
     $x_{j}\leftarrow P^{q}_{u_{j}}(y_{j})$
@@ 4.1 A new bi-level projection
Given a radius $\eta\geq 0$, the goal is to project $Y$ onto the $\ell_{1,\infty}$ norm ball of radius $\eta$, denoted by
@@
In this paper we propose the following alternative new bi-level method. Recall that ${v_{\infty}}=({\|{y_{1}}\|}_{\infty},\dots,{\|{y_{m}}\|}_{\infty})$ is the vector composed of the infinity norms of the columns of matrix $Y$. Let the infinity norm projection $P^{\infty}_{u_{i}}(y)=(\min(y_{i},u_{i}),\forall y_{i}\in y)$. The bi-level $\ell_{1,\infty}$ projection optimization problem is defined by:
$BP^{1,\infty}_{\eta}(Y)=\{X|\forall j,x_{j}=\arg\min\limits_{X_{j}\in\mathcal{B}_{\hat{u}_{j}}^{\infty}}{\|{x_{j}-y_{j}}\|}_{2}$ (13)
@@
  $u\leftarrow P^{1}_{\eta}(({\|{y_{1}}\|}_{\infty},{\|{y_{j}}\|}_{\infty},\dots,{\|{y_{m}}\|}_{\infty}))$
  for $j\in[1,\dots,m]$ do
     $x_{j}\leftarrow P^{\infty}_{u_{j}}(y_{j})$
@@ 4.2 Computational complexity
Synthetic bi-level $\ell_{1,2}$ $\ell_{1,1}$ bi-level $\ell_{1,1}$ $\ell_{1,\infty}$ bi-level $\ell_{1,\infty}$
@@
4.3 Extension to other norms
4.3.1 Extension to the $\ell_{1,1}$ bi-level projection
The $\ell_{1}$ ball is famous for being robust and yielding sparsity. Nevertheless, its extension to matrix, the $\ell_{1,1}$ does not yield structured sparsity. Let ${v_{1}}=({\|{y_{1}}\|}_{1},\dots,{\|{y_{m}}\|}_{1})$ the raw vector composed of the $\ell_{1}$ norm of the columns of matrix $Y$. We propose to define the $\ell_{1,1}$ bi-level optimization problem:
$BP^{1,1}_{\eta}(Y)=\{X|\forall j,x_{j}=\arg\min\limits_{x_{j}\in\mathcal{B}_{\hat{u}_{j}}^{1}}{\|{x_{j}-y_{j}}\|}_{2}$ (14)
@@ 4.3.1 Extension to the $\ell_{1,1}$ bi-level projection
  $u\leftarrow P^{1}_{\eta}(({\|{y_{1}}\|}_{1},\dots,{\|{y_{m}}\|}_{1}))$
  for $j\in[1,\dots,m]$ do
     $x_{j}\leftarrow P_{u_{j}}^{1}(y_{j})$
@@
4.3.2 Extension to bi-level $\ell_{1,2}$ projection
As using $p=1$ yields sparsity in our experiments, we propose to also consider the bi-level $\ell_{1,2}$ projection. Let ${v_{2}}=({\|{y_{1}}\|}_{1},\dots,{\|{y_{m}}\|}_{1})$ the raw vector composed of the $\ell_{2}$ norm of the columns of matrix $Y$. The $\ell_{1,2}$ bi-level optimization problem criterion is:
$BP^{1,2}_{\eta}(Y)=\{X|\forall j,x_{j}=\arg\min\limits_{x_{j}\in\mathcal{B}_{\hat{u}_{i}}^{2}}{\|{x_{j}-y_{j}}\|}_{2}$ (15)
@@ 4.3.2 Extension to bi-level $\ell_{1,2}$ projection
The bi-level projection algorithms for $\ell_{1,2}$ is given by algorithm 4.
@@
  $u\leftarrow P^{1}_{\eta}(({\|{y_{1}}\|}_{2},\dots,{\|{y_{m}}\|}_{2}))$
  for $j\in[1,\dots,m]$ do
     $x_{j}\leftarrow P^{2}_{u_{j}}(y_{j})\forall j\in 1,\dots,m$
@@
Note that the bi-level $\ell_{1,2}$ projection and it’s relationship to other existing norms is out of the scope of this paper. Yet we propose to use this algorithm as a comparison in our experimental section.
@@ 5.1 Tri-level Projection Algorithms
Matrices are not the only mathematical objects that are used in neural networks. Images, for examples, are often represented as order 3 tensors in $\mathbb{R}^{c,n,m}$ with $c$ representing the channels of the image. Moreover, the current development of deep-learning framework for image compression are leaning toward tensor efficient regularization methods [31, 32], and not only vectors or matrices methods. For example, the new image compression standard JPEG AI [33] uses tensor representation of images in the latent space of an autoencoder. That is the reason why we propose to generalize the bi-level projection, first to tri-level, then to multi-level. We consider tensors as multi-dimensional arrays with subscripts indexing. First, consider order 3 tensors of the form $\mathcal{Y}\in\mathbb{R}^{c,n,m}$. For example, $\mathcal{Y}\in\mathbb{R}^{3,m,n}$ is used to represent an image of width $m\times n$ and having 3 channels. We define by $\forall i$ $\in<c,n,m>$ all the tuples of the Cartesian product of the indices. $i$ is a multi-dimensional index, not a single value. For example $\forall i$ <3,256,256$\in$<3,256,256>∈ < 3 , 256 , 256 > $\equiv\forall i\in((1,1,1),(1,1,2),\dots,(3,256,256))$. Given a tensor $\mathcal{Y}\in\mathbb{R}^{c,n,m}$, we denote by $V_{q}=({\|{\mathcal{Y}_{1,1}}\|}_{q},\dots,{\|{\mathcal{Y}_{n,m}}\|}_{q})$ the matrix of aggregation of the channels by the norm $q$.
@@ 6.3.2 Experimental accuracy results
From Table 2 on synthetic dataset and Table 3 on Lung dataset it can be seen that the best accuracy is obtained for $\eta=0.75$ for $\ell_{1,\infty}$ and for $\eta=1.$ for the bilevel $\ell_{1,\infty}$ projection, maximum accuracy of both method are similar. However sparsity and computation time are better for bilevel $\ell_{1,\infty}$ than for regular $\ell_{1,\infty}$.
@@
So bilevel $\ell_{1,\infty}$ projection outperforms usual $\ell_{1,\infty}$ projection in terms of computation cost and sparsity.
@@ 7 Conclusion
Appendix A Iterative form of the multi-level projection